OpenStudy (anonymous):

Write the equation in slope-intercept form, from the given information Hint y=mx+b Passes through (1,2) Slope=7

OpenStudy (anonymous):

y=mx + b given (1,2) and slope(m) = 7 now substitute all of them into the formula to find b. then re-write the whole equation with both the m and b values.

OpenStudy (anonymous):

So: 2=1(7)+b -7 -7 -5=b y=7x-5